OpenStudy (anonymous):

-7(t-4) - (t+ 9) = 5 final answer: 1 3/4???

OpenStudy (anonymous):

-7t + 28 -t -9 = 5 Therefore, -8t +19 = 5 Therefore, 8t = 14 t = 14/8 = 1 3/4 yes. right

OpenStudy (anonymous):

Thanks so much!